OpenStudy (anonymous):

Freddie is at chess practice waiting on his opponent's next move. He notices that the 4-inch-long minute hand is rotating around the clock and marking off time like degrees on a unit circle.

OpenStudy (anonymous):

Are these right? Part 3: How many radians on the unit circle would the minute hand travel from 0° if it were to move 3π inches? l = r(theta) 3pi = 4(theta) 3pi / 4 = theta Part 4: What is the coordinate point associated with this radian measure? \[(\sqrt[-2]{2},\sqrt[2]{2})\] @satellite73 @mathstudent55

OpenStudy (anonymous):

@satellite73 just need someone to check my answers

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@oldrin.bataku

OpenStudy (phi):

the x value is r cos(theta) and the y value is r sin(theta) where theta = 135 deg (or 3 pi/4 radians) and r is 4 inches you get 4 cos(135), 4 sin(135) or \[ ( 4 \cdot -\frac{\sqrt{2}}{2} , 4 \cdot\frac{\sqrt{2}}{2} ) \\ =(-2 \sqrt{2}, \ 2 \sqrt{2}) \] your version has small 2's e.g. \( \sqrt[-2]{2} \) which means something different. You just want to multiply.